# Uncertainty-driven Embedding Convolution

### v2

@@ A.1 Derivation of Bayes-optimal Weights for Uncertainty-driven Embedding Convolution
The optimization problem defined above (Eq. 7) is convex over the probability simplex and has a linear objective. The objective is minimized when more weight is assigned to components with lower values of the cost term $\mathbf{c}_{k}:=\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))+\|\bm{\mu}_{k}(\mathbf{x})\|^{2}$.
We derive the solution using the method of Lagrange multipliers. The Lagrangian is:
$\mathcal{L}(\bm{\pi},\bm{\lambda})=\sum_{k=1}^{K}\bm{\pi}_{k}\mathbf{c}_{k}-\bm{\lambda}\left(\sum_{k=1}^{K}\bm{\pi}_{k}-1\right),$
where $\bm{\lambda}$ is the Lagrange multiplier. Taking the derivative with respect to $\bm{\pi}_{k}$ and setting to zero gives:
$\frac{\partial\mathcal{L}}{\partial\bm{\pi}_{k}}=\mathbf{c}_{k}-\bm{\lambda}=0\quad\Rightarrow\quad\bm{\lambda}=\mathbf{c}_{k}\quad\forall k\ \text{ s.t. }\ \bm{\pi}_{k}>0.$
This condition implies that the minimum is achieved when $\bm{\pi}_{k}$ is inversely proportional to $\mathbf{c}_{k}$. More formally, the optimal solution is:
$\bm{\pi}_{k}^{\star}(\mathbf{x})=\frac{\mathbf{c}_{k}^{-1}}{\sum_{j=1}^{K}\mathbf{c}_{j}^{-1}}=\frac{\left(\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))+\|\bm{\mu}_{k}(\mathbf{x})\|^{2}\right)^{-1}}{\sum_{j=1}^{K}\left(\operatorname{tr}(\bm{\Sigma}_{j}(\mathbf{x}))+\|\bm{\mu}_{j}(\mathbf{x})\|^{2}\right)^{-1}}.$ (8)
In practice, the embeddings $\bm{\mu}_{k}(\mathbf{x})$ are typically $\ell_{2}$-normalized, which implies that $\|\bm{\mu}_{k}(\mathbf{x})\|^{2}\approx 1$ for all $k$. Therefore, we approximate Eq 8 as:
$\bm{\pi}_{k}^{\star}(\mathbf{x})\approx\frac{1/\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))}{\sum_{j=1}^{K}1/\operatorname{tr}(\bm{\Sigma}_{j}(\mathbf{x}))},$
by treating the constant $\|\bm{\mu}_{k}(\mathbf{x})\|^{2}$ term as uniform across all models. This simplification preserves the desirable behavior of assigning higher weights to models with lower predictive uncertainty, while significantly reducing computational overhead.
@@
This optimal solution is Bayes-optimal under the surrogate loss, as it minimizes the expected discrepancy between the ensembled embedding and its positive reference under Gaussian uncertainty. The form of the solution resembles that of precision-weighted model averaging, often seen in Bayesian linear models, where less uncertain models contribute more to the final prediction.
A.2 Derivation of Uncertainty-aware Similarity Estimation
@@
as is common in prior work (Mallik & Sagias, 2011; Randone et al., 2024).
@@
where assume $\mathbf{q}$ and $\mathbf{c}$ are independent.
@@ A.3.1 Preliminary
A.3.2 Approximation of Uncertainty-aware Similarity Estimation
@@
A.3.3 Comparison with squared 2-Wasserstein distance
@@
A.3.4 Comparison with Jeffreys divergence
@@ B.1 Preliminary Analysis of Embedding Subspace Alignment
First, the embedding models considered in our experiments share (i) the same underlying architecture and (ii) largely overlapping training data. As established in prior work (Hewitt & Manning, 2019), such conditions encourage independently trained models to discover consistent linguistic structures—such as syntax, semantic relations, and contextual dependencies—and to encode them in geometrically comparable ways in their embedding spaces.
Second, our ensembling procedure applies $\ell_{2}$ normalization to embeddings before aggregation. This post-processing step has the effect of further mitigating potential subspace differences, effectively aligning embeddings to a common hypersphere and making their scales directly comparable (Timkey & Van Schijndel, 2021).
@@ B.2 Comparison with Alternative Similarity Metrics
In this section, we compare our proposed uncertainty-aware similarity (UEC) against widely used probabilistic similarity metrics. We restrict our comparison to similarity functions that admit closed-form solutions between Gaussian distributions, thereby avoiding the computational overhead of Monte Carlo sampling. Specifically, we consider the KLD (Kullback & Leibler, 1951), 1-Wasserstein (Chhachhi & Teng, 2023), 2-Wasserstein (Gelbrich, 1990), and Closed-form sampled distance (CSD) (Chun, 2023).
@@
We evaluate the similarity metrics under the MIRACL subset retrieval task described in Section 4.1. As shown in Table 7, our proposed UEC achieves the highest performance across all metrics—nDCG@10, Recall@10, and AUC@10—while maintaining the lowest runtime and linear complexity in both memory and time. In contrast, KLD and 2-Wasserstein, while theoretically rich, incur significantly higher computational costs due to matrix inversions and square-root operations, making them less suitable for scalable systems. Notably, UEC does not require any sampling and operates entirely in closed-form, striking an ideal balance between efficiency and uncertainty-awareness. These results underscore the practical advantages of our similarity formulation for real-world deployment.
Table 7: Comparison of similarity metrics under the MIRACL evaluation (Section 4.1). “Rel.” and “Comp.” denote relative runtime (normalized to UEC=1.00) and asymptotic complexity per similarity computation.
Similarity nDCG@10 Recall@10 AUC@10 Memory Complexity Time Rel. Complex. KLD 58.14 78.33 87.60 $\mathcal{O}(KD+KD^{2})$ 1.48 $\mathcal{O}(KD^{2})$ 1‑Wasserstein 57.87 88.21 79.92 $\mathcal{O}(KD)$ 1.21 $\mathcal{O}(KD^{1.5})$ 2‑Wasserstein 57.71 88.48 81.05 $\mathcal{O}(KD+KD^{2})$ 1.60 $\mathcal{O}(KD^{2})$ CSD 58.23 79.10 86.92 $\mathcal{O}(KD)$ 0.88 $\mathcal{O}(KD)$ UEC (Ours) 59.65 80.07 91.04 $\mathcal{O}(KD)$ 1.00 $\mathcal{O}(KD)$
@@ B.3.2 Dataset
For the MIRACL subset evaluation, we constructed a test set by sampling 25 queries each from four languages (English, Russian, Chinese, and Arabic) from the multilingual MIRACL Hard Negative dataset (Zhang et al., 2023). This selection aimed to cover diverse language families and scripts, ensuring a balanced multilingual evaluation.
@@ B.3.4 Hyperparameters
  • UEC (Ours): We fixed the temperature as $1.5$ for all experiments. We search $\beta$ in range of $[0.0001,0.001,0.01,0.1]$.
@@ B.4.4 Hyperparameters
  • UEC (Ours): The temperature was fixed at $1.8$ throughout the experiments. We search $\beta$, scaling parameter for $\sigma_{s}^{2}$, in range of $[0.0001,0.001,0.01,0.1]$.
@@ B.4.5 Additional Tasks
Table 9 summarizes performance on multilingual bitext mining across six representative datasets from MMTEB: Bornholm (Derczynski & Kjeldsen, 2019), Flores (Goyal et al., 2022), IN22Gen (Gala et al., 2023), IndicGenBench (Singh et al., 2024), NTREXB (Federmann et al., 2022), NorwegianCourts (Tiedemann & Thottingal, 2020). For evaluation, we report F1 scores following MMTEB’s standard setup.
Table 9: Comparison of Bitext Mining performance. The best and second-best results per dataset are bolded and underlined. UEC achieves the hightest average performance overall.
@@
The V-Measure is an external clustering evaluation criterion defined as the harmonic mean of homogeneity and completeness (Rosenberg & Hirschberg, 2007). Homogeneity measures whether each cluster contains only members of a single class, while completeness measures whether all members of a given class are assigned to the same cluster. V-Measure ranges from 0 (no alignment) to 100 (perfect clustering), and is widely used in multilingual and scientific document clustering benchmarks such as MMTEB.
@@
Appendix C Statement on the Use of Large Language Models

### v3

@@ A.2 Derivation of Uncertainty-driven Coefficient for Embedding Convolution
The optimization problem in Eq. 10 is strictly convex. Introducing a Lagrange multiplier $\lambda$ for the simplex constraint, the Lagrangian is
$\mathcal{L}(\bm{\pi},\lambda)=\sum_{k}\bm{\pi}_{k}c_{k}+T\sum_{k}\bm{\pi}_{k}\log\bm{\pi}_{k}+\lambda\!\left(\sum_{k}\bm{\pi}_{k}-1\right),$
where $c_{k}:=\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))+\|\bm{\mu}_{k}(\mathbf{x})\|^{2}.$
Setting $\partial\mathcal{L}/\partial\bm{\pi}_{k}=0$ yields
$\log\bm{\pi}_{k}=-\frac{1}{T}c_{k}-\frac{\lambda}{T}-1,$
which implies
$\bm{\pi}_{k}\propto\exp\!\left(-\frac{c_{k}}{T}\right).$
Normalizing over $k$ gives the closed-form solution
$\bm{\pi}_{k}^{\star}(\mathbf{x})=\frac{\exp\!\left(-c_{k}/T\right)}{\sum_{j=1}^{K}\exp\!\left(-c_{j}/T\right)}.$ (11)
Approximation under $\ell_{2}$ Normalization
In practice, embeddings are $\ell_{2}$-normalized, implying $\|\bm{\mu}_{k}(\mathbf{x})\|^{2}\approx 1$ across models. Treating this term as approximately constant, Eq. 11 reduces to
$\bm{\pi}_{k}^{\star}(\mathbf{x})\;\approx\;\frac{\exp\!\left(-\operatorname{tr}(\bm{\Sigma}_{k}(\mathbf{x}))/T\right)}{\sum_{j}\exp\!\left(-\operatorname{tr}(\bm{\Sigma}_{j}(\mathbf{x}))/T\right)},$
which is equivalent to the power-law form in Eq. 4 of the main text.
@@
The resulting coefficients assign higher weight to models with lower epistemic uncertainty for the given query, while entropy regularization ensures smooth, query-adaptive aggregation. This yields a principled uncertainty-aware embedding convolution that is robust to heterogeneous model reliability and distributional shift.
A.3 Derivation of Uncertainty-aware Similarity Estimation
@@
as is common in prior work (Mallik and Sagias, 2011; Randone et al., 2024).
@@
where we assume $\mathbf{q}$ and $\mathbf{c}$ are independent.
@@ A.4.1 Preliminary
A.4.2 Approximation of Uncertainty-aware Similarity Estimation
@@
A.4.3 Comparison with squared 2-Wasserstein distance
@@
A.4.4 Comparison with Jeffreys divergence
@@ B.1 Preliminary Analysis of Embedding Subspace Alignment
First, the embedding models considered in our experiments share (i) the same underlying architecture and (ii) largely overlapping training data. As established in prior work (Hewitt and Manning, 2019), such conditions encourage independently trained models to discover consistent linguistic structures—such as syntax, semantic relations, and contextual dependencies—and to encode them in geometrically comparable ways in their embedding spaces.
Second, our ensembling procedure applies $\ell_{2}$ normalization to embeddings before aggregation. This post-processing step has the effect of further mitigating potential subspace differences, effectively aligning embeddings to a common hypersphere and making their scales directly comparable (Timkey and Van Schijndel, 2021).
@@ B.2 Comparison with Alternative Similarity Metrics
In this section, we compare our proposed uncertainty-aware similarity (UEC) against widely used probabilistic similarity metrics. We restrict our comparison to similarity functions that admit closed-form solutions between Gaussian distributions, thereby avoiding the computational overhead of Monte Carlo sampling. Specifically, we consider the KLD (Kullback and Leibler, 1951), 1-Wasserstein (Chhachhi and Teng, 2023), 2-Wasserstein (Gelbrich, 1990), and Closed-form sampled distance (CSD) (Chun, 2023).
@@
We evaluate the similarity metrics under the MIRACL subset retrieval task described in Section 4.1. As shown in Table 9, UEC achieves the best overall retrieval performance while exhibiting the lower runtime.
We note that the asymptotic complexities of KLD and Wasserstein distances can be reduced to $\mathcal{O}(KD)$ when restricted to diagonal covariance matrices. However, despite having the same big-$\mathcal{O}$ complexity, these metrics remain consistently slower than UEC in practice. This gap arises from fundamental differences in their computational structure. Specifically, KLD and Wasserstein still require per-dimension normalization, division, logarithmic, or square-root operations, as well as multiple intermediate tensor constructions. In contrast, UEC consists solely of element-wise additions and multiplications followed by a single aggregation, with no normalization constants, matrix operations, or transcendental functions.
As a result, UEC has a substantially smaller constant factor and lower memory access overhead, enabling faster execution and better hardware utilization, particularly in large-scale retrieval settings. Moreover, UEC operates entirely in closed form without sampling or iterative solvers, making it well-suited for deployment in latency-sensitive systems.
Table 9: Comparison of similarity metrics under the MIRACL evaluation (Section 4.1). “Rel.” and “Comp.” denote relative runtime (normalized to UEC=1.00) and asymptotic complexity per similarity computation. Although several metrics share the same asymptotic complexity under diagonal covariance assumptions, their empirical runtimes differ due to operation-level costs and implementation overheads, which are not captured by big-$\mathcal{O}$ notation.
Similarity nDCG@10 Recall@10 AUC@10 Memory Complexity Time Rel. Complex. KLD 58.14 78.33 87.60 $\mathcal{O}(KD)$ 1.52 $\mathcal{O}(KD)$ 1-Wasserstein 57.87 88.21 79.92 $\mathcal{O}(KD)$ 1.21 $\mathcal{O}(KD)$ 2-Wasserstein 57.71 88.48 81.05 $\mathcal{O}(KD)$ 2.37 $\mathcal{O}(KD)$ CSD 58.23 79.10 86.92 $\mathcal{O}(KD)$ 0.88 $\mathcal{O}(KD)$ UEC (Ours) 59.65 80.07 91.04 $\mathcal{O}(KD)$ 1.00 $\mathcal{O}(KD)$
@@ B.3.2 Dataset
For the MIRACL subset evaluation, we constructed a test set by sampling 100 queries each from four languages (English, Russian, Chinese, and Arabic) from the multilingual MIRACL Hard Negative dataset (Zhang et al., 2023). This selection aimed to cover diverse language families and scripts, ensuring a balanced multilingual evaluation.
@@ B.3.4 Hyperparameters
  • UEC (Ours): We fixed the temperature $T$ as $1.5$ for all experiments. We search $\beta$ in range of $[0.0001,0.001,0.01,0.1]$.
@@ B.4.4 Hyperparameters
  • UEC (Ours): The temperature $T$ was fixed at $1.8$ throughout the experiments. We search $\beta$, scaling parameter for $\sigma_{s}^{2}$, in range of $[0.0001,0.001,0.01,0.1]$.
@@ B.4.5 Additional Tasks
Table 11 summarizes performance on multilingual bitext mining across six representative datasets from MMTEB: Bornholm (Derczynski and Kjeldsen, 2019), Flores (Goyal et al., 2022), IN22Gen (Gala et al., 2023), IndicGenBench (Singh et al., 2024), NTREXB (Federmann et al., 2022), NorwegianCourts (Tiedemann and Thottingal, 2020). For evaluation, we report F1 scores following MMTEB’s standard setup.
Table 11: Comparison of Bitext Mining performance. The best and second-best results per dataset are bolded and underlined. UEC achieves the highest average performance overall.
@@
The V-Measure is an external clustering evaluation criterion defined as the harmonic mean of homogeneity and completeness (Rosenberg and Hirschberg, 2007). Homogeneity measures whether each cluster contains only members of a single class, while completeness measures whether all members of a given class are assigned to the same cluster. V-Measure ranges from 0 (no alignment) to 100 (perfect clustering), and is widely used in multilingual and scientific document clustering benchmarks such as MMTEB.
@@
B.5 Multilingual Sets
Retrieval
In multilingual retrieval, UEC consistently outperforms all baselines. Across seven languages, UEC attains the best average nDCG@10, Recall@100, and AUC@10, while never degrading performance on any metric compared to the strongest individual model. Notably, UEC yields substantial gains on low-resource languages such as Bengali and Hindi, where single models exhibit complementary strengths: BGE and E5 provide strong general-purpose multilingual retrieval, whereas GTE-MB is tailored to multilingual benchmarks. By reweighting models according to their epistemic uncertainty, UEC effectively selects the most reliable expert per query, improving both top-ranked retrieval quality and the calibration of uncertainty curves (AUC@10).
Table 14: Comparison of retrieval performance on the WikipediaRetrievalMultilingual benchmark. The best and second-best results per row are bolded and underlined. UEC achieves the highest average performance across all languages and metrics, with particularly strong gains on AUC@10, indicating more reliable uncertainty-aware ranking. (Uni.=Uniform, Wt.=Weighted, TA=Task arithmetic.).
Individual Models Model Merging Ensemble Language / Metric BGE E5 GTE GTE-MB Uni. Wt. TA Uni. Wt. UEC Bulgarian     nDCG@10 29.11 45.61 33.97 55.29 35.62 39.12 38.87 32.51 38.88 48.03     Recall@100 52.86 71.60 58.13 81.53 63.40 68.39 68.24 56.33 67.63 74.12     AUC@10 57.36 62.63 55.06 62.12 58.29 60.16 59.93 56.32 60.77 65.28 Bengali     nDCG@10 7.33 8.11 5.54 3.16 6.52 7.83 7.74 6.36 7.91 9.86     Recall@100 21.06 23.93 17.40 10.46 19.75 21.97 21.03 19.66 23.74 25.73     AUC@10 39.71 19.28 29.58 47.36 26.83 24.97 36.22 31.65 30.78 45.97 Danish     nDCG@10 63.45 75.18 71.16 69.91 67.83 73.42 72.11 68.94 75.14 76.86     Recall@100 87.26 94.33 92.00 91.33 91.24 92.47 92.33 91.49 93.81 95.03     AUC@10 64.22 73.07 65.35 68.92 66.67 69.39 69.72 68.04 73.01 76.28 Persian     nDCG@10 9.19 14.44 8.97 35.69 11.36 13.28 12.97 12.06 14.23 16.92     Recall@100 26.80 33.40 27.73 61.33 28.83 30.05 30.16 29.97 33.32 37.73     AUC@10 33.55 39.01 28.95 59.47 34.26 36.83 35.55 35.92 39.23 45.19 Hindi     nDCG@10 11.99 21.55 5.23 17.16 15.37 20.05 17.83 16.32 21.11 23.56     Recall@100 29.33 45.93 14.87 31.33 31.62 42.18 40.17 39.20 45.02 47.22     AUC@10 37.84 49.91 46.51 53.29 46.97 48.64 48.03 50.26 49.99 52.38 Portuguese     nDCG@10 72.29 81.12 75.40 75.94 75.32 79.17 78.93 76.66 80.64 83.13     Recall@100 92.73 96.13 95.13 95.66 94.83 95.27 95.21 95.17 96.08 96.94     AUC@10 54.28 71.07 68.13 70.52 64.82 70.34 68.73 71.26 72.35 77.17 Serbian     nDCG@10 31.95 41.52 33.99 44.15 35.52 41.19 41.82 37.12 41.89 44.38     Recall@100 54.20 67.73 58.40 74.06 64.93 67.71 69.74 69.18 72.09 74.81     AUC@10 21.25 63.42 67.68 61.25 54.97 63.02 64.11 62.65 66.16 69.21 Avg. nDCG@10 $\uparrow$ 32.18 41.07 33.47 43.04 35.36 39.15 38.61 35.71 39.97 43.24 Avg. Recall@100 $\uparrow$ 52.03 61.86 51.95 63.52 56.37 59.72 59.55 57.28 61.67 64.51 Avg. AUC@10 $\uparrow$ 44.03 54.05 51.61 60.41 50.40 53.34 54.61 53.73 56.05 61.64
Classification
On SwissJudgementClassification, UEC yields the best average Accuracy, F1, and AUROC across all three languages. While GTE-MB is competitive as a strong multilingual encoder, it is dominated by UEC once epistemic uncertainty is taken into account at the ensemble level. Crucially, UEC does not trade off performance between language groups: it matches or improves upon the strongest baseline in German and French, and achieves the largest gains on Italian, the smallest and most imbalanced language split. This suggests that UEC remains robust under multilingual label imbalance and does not sacrifice minority-language performance when down-weighting uncertain embeddings.
Table 15: Comparison of multilingual legal judgment classification performance on SwissJudgementClassification. The best and second-best results per row are bolded and underlined. UEC achieves the highest average scores over German, French, and Italian, demonstrating robust performance across languages and label distributions. (Uni.=Uniform, Wt.=Weighted, TA=Task arithmetic.)
Individual Models Model Merging Ensemble Language / Metric BGE E5 GTE GTE-MB Uni. Wt. TA Uni. Wt. UEC German     Accuracy 50.28 47.85 47.73 53.08 48.26 49.73 48.08 49.14 51.75 53.13     F1 43.86 41.96 43.01 46.10 42.67 44.61 44.43 43.24 44.61 45.97     AUROC 50.32 50.17 50.86 51.12 50.17 50.21 50.83 50.72 50.26 51.83 French     Accuracy 57.16 58.58 56.37 55.45 56.82 55.90 55.44 57.09 57.71 58.36     F1 46.63 47.25 46.34 46.41 46.55 45.87 45.73 46.61 47.02 48.52     AUROC 50.48 50.22 50.65 50.67 50.13 50.24 50.37 50.53 50.51 51.02 Italian     Accuracy 53.47 57.61 53.00 58.36 54.76 55.89 54.83 54.92 57.65 58.72     F1 45.64 46.09 45.72 47.79 45.88 46.03 45.31 45.62 46.72 48.03     AUROC 50.73 50.39 50.42 51.36 50.89 51.25 51.08 50.56 50.84 51.89 Avg. Accuracy $\uparrow$ 53.64 54.68 52.37 55.63 53.28 53.84 52.79 53.72 55.65 56.74 Avg. F1 $\uparrow$ 47.38 45.10 45.02 46.77 45.03 45.50 45.16 45.16 46.12 47.51 Avg. AUROC $\uparrow$ 50.51 50.26 50.64 51.05 50.40 50.57 50.76 50.60 50.54 51.58
STS
On SemRel24STS, UEC further improves over strong multilingual encoders on sentence-level similarity. While GTE-MB and E5 already achieve competitive performance on high-resource languages such as English and Afrikaans, UEC attains the best average score and ranks first on 9 out of 12 languages, including several low-resource African and Indic languages (e.g., Hausa, Hindi, Marathi). This indicates that uncertainty-driven ensembling not only preserves the strengths of specialist models but also yields more stable semantic similarity estimates in challenging multilingual and low-resource regimes.
Table 16: Comparison of multilingual STS performance on SemRel24STS. The best and second-best results per column are bolded and underlined. UEC attains the highest average Spearman correlation, ranking first on 9 out of 12 languages and consistently outperforming both individual encoders and heuristic ensembles. We evaluate on 12 languages: Afrikaans (Afr.), Amharic (Amh.), Modern Standard Arabic (Arb.), Algerian Arabic (Arq.), Moroccan Arabic (Ary.), English (Eng.), Hausa (Hau.), Hindi (Hin.), Indonesian (Ind.), Kinyarwanda (Kin.), Marathi (Mar.), and Telugu (Tel.). (Uni.=Uniform, Wt.=Weighted, TA=Task arithmetic.).
Model / Language Afr. Amh. Arb. Arq. Ary. Eng. Hau. Hin. Ind. Kin. Mar. Tel. Avg. Individual BGE 72.13 12.28 30.88 37.80 22.93 79.52 33.69 48.91 46.99 45.26 55.00 27.43 42.73 E5 77.16 15.43 27.58 40.42 12.35 81.37 41.30 47.28 51.99 49.41 54.20 29.25 43.97 GTE 77.29 15.21 22.01 34.19 12.88 77.79 35.25 40.69 45.20 47.33 45.66 32.36 40.48 GTE-MB 77.03 18.95 28.42 42.34 31.32 75.02 43.70 24.11 42.71 49.95 44.30 35.81 42.81 Model Merging Uni. 75.87 13.43 26.91 38.24 14.18 78.99 38.42 44.34 48.09 47.01 52.31 28.87 42.22 Wt. 76.34 14.17 27.94 38.43 13.98 80.05 38.79 46.26 49.73 48.28 52.44 28.49 42.91 TA 76.21 13.78 27.33 38.32 14.08 79.43 38.51 45.58 49.22 47.49 52.33 28.54 42.57 Ensemble Uni. 75.79 13.82 27.18 38.19 14.25 79.16 38.19 44.85 48.34 46.91 52.42 28.99 42.34 Wt. 76.46 15.12 28.06 39.18 13.83 80.86 39.44 46.52 50.93 48.88 53.02 28.26 43.38 UEC 77.84 16.77 31.02 40.88 23.88 82.28 43.81 50.11 52.46 50.56 55.48 34.17 46.61
Appendix C Uncertainty Estimation Diagnosis
C.1 Uncertainty Estimation of Post-hoc Probabilistic Embedding Model
We evaluate the behavior of the Laplace-based post-hoc probabilistic embedding model under several factors: covariance structure, the number of layers included in the curvature approximation, and the amount of data used for fitting. For all experiments, we fit LA following the same configuration in Appendix B.4, and evaluate uncertainty using a binary classification proxy task on MS MARCO.
Covariance Structure
We first compare diagonal and KFAC covariance structures applied to the same penultimate-layer parameters with E5 model. KFAC (Kronecker-Factored Approximate Curvature) approximates the full Fisher or Hessian matrix using a Kronecker factorization of layerwise curvature, enabling structured (second-order) uncertainty modeling that is substantially richer than a diagonal approximation. In principle, this allows correlations between parameters within the same linear layer to be captured, providing a more expressive posterior.
Table 17: Diagonal vs. KFAC covariance for LA. $m$, $d$, and $N$ denote penultimate-layer width, embedding dimension, and number of LA fitting samples, respectively. Diagonal covariance achieved near identical calibration performance without extra computational overhead compared to KFAC.
Cov Type ACC $\uparrow$ ECE $\downarrow$ Time Comp. Mem. Comp.
Diag 83.57 0.036 $\mathcal{O}(Nmd)$ $\mathcal{O}(md)$
KFAC 83.57 0.028 $\mathcal{O}\!\left(N(m^{2}{+}d^{2})+m^{3}{+}d^{3}\right)$ $\mathcal{O}(m^{2}{+}d^{2})$
Table 17 shows that both predictive accuracy and calibration (as measured by ECE) are nearly identical across the two settings. Despite capturing a higher-fidelity curvature structure, the KFAC variant does not yield measurable improvements in uncertainty quality for this setting while incurring significantly larger computational overhead. Given this trade-off, we adopt the diagonal covariance in the main experiments.
Number of Layers
To examine whether deeper curvature modeling affects post-hoc uncertainty, we vary the number of layers included in the LA from only the final linear projection layer (the standard configuration) up to six layers.
Table 18: Effect of increasing the number of layers included in LA. Extending the curvature approximation to multiple layers yields negligible gains while significantly increasing computational cost.
# Layers 1 (Last) 2 4 6
ACC $\uparrow$ 83.57 83.57 83.56 83.57
ECE $\downarrow$ 0.036 0.035 0.038 0.035
As shown in Table 18, both accuracy and ECE remain stable across all configurations. This behavior aligns with recent findings in Daxberger et al. (2021) and Sharma et al. (2023), which show that last-layer LA already provide strong and computationally efficient uncertainty estimates for large pretrained encoders. Our observations support this conclusion: extending the curvature approximation to multiple layers yields negligible gains while significantly increasing computational cost.
Data Sparsity
We further examine how the amount of data used for fitting the Laplace Approximation affects uncertainty estimation. We subsample the fitting set to $50\%$, $10\%$, and $1\%$ of the original 7.5K samples, and additionally include a $0\%$ condition corresponding to purely deterministic embeddings without LA to isolate the contribution of the posterior variance.
Table 19: Effect of Laplace fitting data sparsity on uncertainty estimation. Calibration error increases gradually under extreme sparsity, but remains well-behaved for moderate subsampling.
Sparsity 100% (Original) 50% 10% 1% 0% (Det.)
ACC $\uparrow$ 83.57 83.57 83.56 83.56 83.57
ECE $\downarrow$ 0.036 0.038 0.051 0.061 0.063
Table 19 shows that accuracy is unaffected by data sparsity, confirming that LA does not harm predictive quality. More importantly, all LA–based settings outperform the deterministic baseline ($0\%$) in calibration. Even with only $1\%$ of the fitting data ($\sim$75 samples), LA yields mild but consistent calibration gains. With $10\%$ of the data ($\sim$750 samples), the calibration error nearly matches the full-data setting, indicating that LA requires only a modest amount of data to produce stable variance estimates. Overall, these results demonstrate that LA is not particularly sensitive to noise or sparsity, and that uncertainty estimation remains reliable even under substantial subsampling.
C.2 Variance Calibration of $\sigma_{s}^{2}$
To directly assess the quality of $\sigma_{s}^{2}$ as a variance estimator, we perform a variance–error calibration analysis, as mentioned in Section 4.3.2.
Setup
Standard calibration metrics such as ECE are designed for probability outputs and are not directly applicable to variance-based confidence measures such as $\sigma_{s}^{2}$. To address this limitation, we introduce Var-ECE, a variance-oriented calibration metric that evaluates whether lower predicted variance corresponds to higher empirical correctness. For each sentence pair $(x_{1},x_{2})$ in SemRel24STS, a human similarity score $y\in[0,1]$ is provided. Under UEC, the similarity is modeled as
$s(x_{1},x_{2})\sim\mathcal{N}\big(\mu_{s}(x_{1},x_{2}),\,\sigma_{s}^{2}(x_{1},x_{2})\big),$
where $\mu_{s}$ is the predicted mean similarity and $\sigma_{s}^{2}$ is the epistemic variance. If $\sigma_{s}^{2}$ is an accurate estimate of predictive variance, then examples with large predicted variance should on average incur larger squared error. We therefore compute the empirical squared error
$r(x_{1},x_{2})\;\coloneqq\;\big(\mu_{s}(x_{1},x_{2})-y(x_{1},x_{2})\big)^{2},$
and directly compare $\sigma_{s}^{2}$ and $r$ across groups of examples.
Variance–Error Binning
To reduce noise and obtain stable estimates, we partition the predicted variances into $M$ quantile bins $\{B_{m}\}_{m=1}^{M}$ based on $\sigma_{s}^{2}(x_{1},x_{2})$ (we use $M=10$ in all experiments). For each bin, we compute the average predicted variance and the average squared error:
$\overline{\sigma^{2}}_{m}=\frac{1}{|B_{m}|}\sum_{(x_{1},x_{2})\in B_{m}}\sigma_{s}^{2}(x_{1},x_{2}),\qquad\overline{r}_{m}=\frac{1}{|B_{m}|}\sum_{(x_{1},x_{2})\in B_{m}}r(x_{1},x_{2}).$
If $\sigma_{s}^{2}$ is well estimated, then $\overline{\sigma^{2}}_{m}$ should be close to $\overline{r}_{m}$ for every bin: examples assigned higher predicted variance should indeed exhibit higher empirical error.
Variance Calibration Error (Var-ECE)
We measure the discrepancy between predicted and empirical variances via
$\text{Var-ECE}\;=\;\sum_{m=1}^{M}\frac{|B_{m}|}{N}\left|\ \overline{r}_{m}-\overline{\sigma^{2}}_{m}\right|,$ (31)
where $N$ is the total number of evaluated examples. A small Var-ECE indicates that the predicted variance $\sigma_{s}^{2}$ numerically matches the observed squared error across uncertainty levels, meaning the UEC provides an accurate estimate of predictive variance.
Appendix D Sensitivity to $\beta$ and Temperature $T$
We analyze the sensitivity of UEC to two key hyperparameters: (1) the coefficient $\beta$ that determines how strongly the estimated variance influences the uncertainty-aware similarity, and (2) the temperature parameter $T$ that sharpens the ensemble coefficients.
MIRACL Subset
Figure 6 summarizes the effect of varying $\beta$ and $T$ on MIRACL Subset. While the curves exhibit noticeable movement, UEC maintains stable behavior across a broad range of values. This reflects that both hyperparameters meaningfully control how uncertainty modulates similarity and coefficient sharpening. Performance consistently peaks near $\beta=10^{-2}$ and $T=1.5$, which provides a balanced trade-off across all retrieval metrics.
(a) Sensitivity to $\beta$
(b) Sensitivity to temperature $T$
Figure 6: Sensitivity analysis of UEC on MIRACL. Both $\beta$ and $T$ influence performance in interpretable ways controlling variance contribution and sharpening, respectively. UEC exhibits a stable optimum region around $(\beta=10^{-2},\,T=1.5)$, making tuning straightforward in practice.
MMTEB
A similar analysis on the StackOverflow task (MMTEB) is presented in Figure 7. Remarkably, the optimal region remains highly consistent with the MIRACL case: the best performance again emerges around $\beta=10^{-2}$ and requires only mild temperature sharpening (here, $T\approx 1.8$).
(a) Sensitivity to $\beta$
(b) Sensitivity to temperature $T$
Figure 7: Sensitivity analysis of UEC on StackOverflow (MMTEB). Despite dataset differences, the optimal hyperparameter region closely matches that of MIRACL, indicating that UEC’s hyperparameter tuning is stable and transferable across tasks.
Appendix E Statement on the Use of Large Language Models